# Semantic Web: Past, Present, and Future

### v1

@@ 4.2 Ontologies
An ontology is commonly defined as a formal, machine-readable representation of key concepts and relationships within a specific domain [122, 120]. In essence, ontologies capture a shared perspective [122] that is, the formal conceptualization of ontologies expresses a consensus view among different stakeholders. Visualizing ontologies is akin to viewing a spectrum, with a specificity of concepts, their relationships, and the granularity of meaning representation varying along this continuum  [109, 160, 159]. A controlled vocabulary corresponds to the less expressive form of ontology, comprising a restrictive list of words or terms used for labeling, indexing, or categorization. The Clinical Data Interchange Standards Consortium (CDISC) Terminology is an exemplary vocabulary that harmonizes definitions in clinical research. [^42] https://datascience.cancer.gov/resources/cancer-vocabulary/cdisc-terminology A thesaurus is located next in the spectrum; they enhance controlled vocabularies with information about terms and their synonyms and broader/narrower relationships. The Unified Medical Language System (UMLS) integrates medical terms and their synonyms. [^43] https://www.nlm.nih.gov/research/umls/index.html Next, taxonomies are built over controlled vocabularies to provide a hierarchical structure, e. g., parent/child relationship. SNOMED-CT [^44] https://www.snomed.org/value-of-snomedct (Systematized Nomenclature of Medicine Clinical Terms) provides a terminology and coding system used in healthcare and medical fields; medical concepts organized in a hierarchical structure enabling a granular representation of clinical information. The Simple Knowledge Organization System (SKOS) [^45] https://www.w3.org/TR/skos-reference/ is a W3C standard to describe knowledge about organizational systems. Lastly, ontologies are at the highest extreme of the spectrum, integrating sets of concepts with attributes and relationships to define a domain of knowledge.
Note, SKOS is a popular standard for modeling domain-specific taxonomies in the different scientific communities such as economics, social sciences, etc. to represent concepts and their relationships, most importantly narrower, broader, and related. However, it does not have the expressiveness of OWL with its complex expressions on classes and relations. For a detailed discussion, we refer to the literature such as [96] and the W3C on using OWL and SKOS [^46] https://www.w3.org/2006/07/SWD/SKOS/skos-and-owl/master.html.
@@ 4.3 Types and Examples of Ontologies
A network of ontologies, such as the example shown in Figure 1, may consist of a variety of ontologies created by different actors and communities. Ontologies may be the result of a transformation or reengineering activity of a legacy system, such as a relational database or existing taxonomy such as the Dewey Decimal Classification [^47] http://dewey.info/ or Dublin Core. Other ontologies are created from scratch. This involves applying existing methods and tools for ontology engineering and choosing an appropriate representation language for the ontology (see Section 6).
@@
Ontology engineering deals with the methods for creating ontologies [70] and has its origins in software engineering in the creation of domain models and in database design in the creation of conceptual models. A good overview of ontology engineering can be found in several reference books [70]. Ontologies vary greatly in their structure, size, development methods applied, and application domains considered. Complex ontologies are also distinguished in terms of their purpose and granularity.
@@
Domain Ontologies represent knowledge specific to a particular domain [53, 120]. Domain ontologies are used as external sources of background knowledge [53]. They can be built on foundational ontologies [121] or core ontologies [142], which provide precise structuring to the domain ontology and thus improve interoperability between different domain ontologies. Domain ontologies can be simple such as the FOAF ontology or the event ontology mentioned above, or very complex and extensive, having been developed by domain experts, such as the SNOMED medical ontology.
@@
Core Ontologies represent a precise definition of structured knowledge in a particular domain spanning multiple application domains [142, 120]. Examples of core ontologies include core ontologies for software components and web services [120], for events and event relationships [140], or for multimedia metadata [137]. Core ontologies should thereby build on foundational ontologies to benefit from their formalization and strong axiomatization [142]. For this purpose, new concepts and relations are added to core ontologies for the application domain under consideration and are specialized by foundational ontologies.
@@
Foundational Ontologies have a very wide scope and can be reused in a wide variety of modeling scenarios [27]. They are therefore used for reference purposes [120] and aim to model the most general and generic concepts and relations that can be used to describe almost any aspect of our world [27, 120], such as objects and events. An example is the Descriptive Ontology for Linguistic and Cognitive Engineering (DOLCE) [27]. Such basic ontologies have a rich axiomatization that is important at the developmental stage of ontologies. They help ontology engineers to have a formal and internally consistent conceptualization of the world, which can be modeled and checked for consistency. For the use of foundational ontologies in a concrete application, i. e., during the runtime of an application, the rich axiomatization can often be removed and replaced by a more lightweight version of the foundational ontology.
@@ 4.4 Distributed Network of Ontologies and Ontology Patterns
A network of ontologies must be flexible with respect to the functional requirements imposed on it. This is because systems are modified, extended, combined, or integrated over time. In addition, the networked ontologies must lead to a common understanding of the modeled domain. This common understanding can be achieved through a sufficient level of formalization and axiomatization, and through the use of ontology patterns. An ontology pattern, similar to a design pattern in software engineering, represents a generic solution to a recurring modeling problem [142]. Ontology patterns allow to select parts from the original ontology. Either all or only certain patterns of an ontology can be reused in the network. Thus, to create a network of ontologies, e. g., existing ontologies and ontology patterns can be merged on the Web. The ontology engineer can drive or explicitly provide for the modularization of ontologies using ontology patterns. Core ontologies represent one approach to designing a network of ontologies (see in detail [142]). They allow to capture and exchange structured knowledge in complex domains. Well-defined core ontologies fulfill the properties mentioned in the previous section and allow easy integration and smooth interaction of ontologies (see also [142]). The networked ontologies approach leads to a flat structure, as shown in Figure 1, where all ontologies are used on the same level. Such structures can be managed up to a certain level of complexity.
@@
The approach of networked core ontologies is illustrated by the example of ontology layers starting from foundational to core to domain ontologies. As shown in Figure 4, DOLCE is the foundational ontology at the bottom layer, the Multimedia Metadata Ontology (M3O) [137] as the core ontology for multimedia metadata, and an extension of M3O for the music domain. Core ontologies are typically defined in description logic and cover a field larger than the specific application domain requires [62]. Concrete information systems will typically use only a subset of core ontologies. To achieve modularization of core ontologies, they should be designed using ontology patterns. By precisely matching the concepts in the core ontology with the concepts provided in the foundational ontology, they provide a solid foundation for future extensions. New patterns can be added and existing patterns can be extended by specializing the concepts and roles. Figure 4 shows different patterns of the M3O and DOLCE ontologies.
Ideally, the ontology patterns of the core ontologies are reused in the domain ontologies [62], as shown in Figure 4. However, since it cannot be assumed that all domain ontologies are aligned with a foundational or core ontology, the option that domain ontologies are developed and maintained independently must also be considered. In this case, domain knowledge can be reused in core ontologies by applying the Descriptions and Situations (DnS) ontology pattern of the foundational ontology DOLCE. The DnS ontology pattern is an ontological formalization of context [120] by defining different views using roles. These roles can refer to domain ontologies and allow a clear separation of the structured knowledge of the core ontology and domain-specific knowledge. To model a network of ontologies, such as the example described above, the Web Ontology Language (OWL) and its ability to axiomatize using description logic [15] is used. In addition to being used to model a distributed knowledge representation and integration, OWL, is also used in particular to derive inferences from this knowledge, which is described in Section 6.
@@ 5.1 Graph Data Creation
Graph data can be created by transforming legacy data via a data integration system [106], which consists of a unified schema, data sources, and mapping rules. These mapping rules define the concepts within the schema and establish links to the data sources. By employing declarative definitions, knowledge graph creation promotes modularity and reusability. This approach allows users to trace the entire graph creation process, leading to improved transparency and ease of maintenance.
@@
The mapping language R2RML [41] defines mapping rules from relational databases (relational data models) to RDF graphs. These mappings themselves are also RDF triples [18]. Because of its compact representation, Turtle is considered a user-friendly notation of RDF graphs. The structure of R2RML is illustrated in Figure 5; essentially, table contents are mapped to triples by the classes SubjectMap, PredicateMap, and ObjectMap. If the object is a reference to another table, this reference is called RefObjectMap. Here, SubjectMap contains primary key attributes of the corresponding table. Thus, there exists a mapping rule representable in RDF graphs by means of which tables of relational databases can be represented as RDF graphs.
Figure 5: Structure of a relational data mapping (source: [41]).
@@
The RDF Mapping Language (RML)[46] extends R2RML to encompass the definition of logical sources in various formats, including CSV, JSON, XML, and HTML. This enhancement enables RML to introduce new operators that facilitate the integration of data from diverse sources into the Semantic Web. Thus, instead of LogicalTable, RML includes the tag LogicalSource, to allow for the retrieval of data in several formats. Additionally, RML resorts to W3C-standardized vocabularies and enables the definition of retrieval procedures to collect data from Web APIs or databases. R2RML and RDF mapping rules are expressed in RDF, and their graphs document how classes and properties in one or various ontologies that are part of an RDF graph are populated from data collected from potentially heterogeneous data sources.
@@
Over time, the Semantic Web community has actively contributed to addressing the challenge of integrating heterogeneous datasets, resulting in the development of several frameworks for executing declarative mapping rules [37, 31, 127]. A rich spectrum of tools (e. g., RMLMapper [46], RocketRML [147], CARML [^48] https://github.com/carml/carml, SDM-RDFizer [94], Morph-KGC [9], and RMLStreamer [123]) offers the possibility of executing R2RML and RML rules and efficiently materializing the transformed data into RDF graphs. Van Assche et al. [11] provided an extensive survey detailing the main characteristics of these engines. Despite significant efforts in developing these solutions, certain parameters can impact the performance of the graph creation process [34]. Existing engines may face challenges when handling complex mapping rules or large data sources. Nonetheless, the community continues to collaborate and address these issues. An example of such collaboration is the Knowledge Graph Construction Workshop 2023 Challenge [^49] https://zenodo.org/record/7689310 that took place at ESWC 2023. This community event aims to understand the strengths and weaknesses of existing approaches and devise effective methods to overcome existing limitations.
@@
RDF graphs can also be dynamically created through the execution of queries over data sources. These queries involve the rewriting of queries expressed in terms of an ontology, based on mapping rules that establish correspondences between data sources and the ontology. Tools such as Ontop [31], Ultrawrap [146], Morph [127], Squerall [108], and Morph-CSV [35] exemplify systems that facilitate the virtual creation of RDF graphs.
@@ 5.2 Quality and Validation of Graph Data
Quality and validation of the graph data are crucial to maintaining the integrity of the Semantic Web [48, 42, 173]. The evaluation of integrity constraints allows for the identification of inconsistencies, inaccuracies, or contradictions within the data. They also help maintain consistency by ensuring related data elements remain coherent. Constraints are logical statements – expressed in a particular language – that impose restrictions on the values taken for target nodes in a given property.
Constraints can be expressed using OWL [155], SPARQL queries [105], or using shapes. However, the interpretation of the results depends on the semantics followed to interpret the failure of an integrity constraint. For example, constraints expressed in OWL are validated using an Open-World Assumption (OWA) (i. e., a statement cannot be inferred to be false based on failures to prove it) and under the absence of the Unique Name Assumption (UNA) (i. e., two different names may refer to the same object). These two features make it difficult to validate data in applications where data is supposed to be complete. Definitions of integrity constraint semantics in OWL using the Closed-World Assumption [114, 115, 155] overcome these issues.
@@
Data quality conditions and integrity constraints can also be expressed as graphs of shapes or the so-called shapes schema. A shape corresponds to a conjunction of constraints that a set of nodes in an RDF graph must satisfy [86]. These constraints can restrict the types of nodes, the cardinality of certain properties, and the expected data types or values for specific properties. A shape can target the instances of a class, the domain or range of a property, or a specific node in the RDF graph. A shape or node in a shapes graph is validated in an RDF graph, if and only if, all the target nodes in the RDF graph satisfy all the constraints in the shape. Figure 6 presents a shapes graph of three shapes targeting the classes Brand, Playcount, and MusicArtist. Each of the shapes comprises one constraint. In the shapes Brand and MusicArtist the properties title and name can take more than one value, while the shape Playcount states that each instance of the class Playcount must have exactly one value of the property count. Additionally, the instances of the class Brand must be related to valid instances of the class Playcount which should also be related to valid instances of the class MusicArtist.
@@
There are two standards for defining shapes, ShEx (Shape Expressions) [128]) and SHACL (Shapes Constraint Language) [103]). Both define shapes over the attributes (i. e., owl:DatatypeProperties), and constraints on incoming/outgoing arcs, cardinalities, RDF syntax, and extension mechanism. These inter-class constraints induce a shape network used to validate the integrity and data quality properties of an RDF graph.
@@
SHACL and ShEx, although sharing a common goal, adopt distinct approaches. ShEx seeks to offer a language serving as a grammar or schema for RDF graphs, delineating RDF graph structures for validation. On the other hand, SHACL is positioned as the W3C recommendation for validating RDF graphs against a conjunction of constraints, emphasizing a constraint language for RDF. Despite their analogous roles in specifying shapes and constraints for RDF data, ShEx and SHACL differ in syntax, expressiveness, and community adoption [64].
@@
The evaluation results of a SHACL shape network over an RDF graph are presented in validation reports using a controlled vocabulary. A validation report includes explanations about the violations, the severity of the violation, and a message describing the violation. SHACL is the language selected by the International Data Space (IDS) to express the restrictions that state the integrity over RDF graphs [104]. Besides the integrity validation of an RDF graph, SHACL can be utilized to describe data sources and the certification of a query answer [135], as metadata to enhance the performance of a SPARQL query engine [129], to certify access policies [136], and to provide provenance as a result of the validation of integrity constraints [44].
@@
In the context of a quality assessment pipeline, one crucial step involves validating the shape schema against a graph. It is important to mention that the validation of recursive shape schemas is not explicitly addressed in the SHACL specification [103]. To address this gap, Corman et al. [40] introduce a semantic framework for validating recursive SHACL. They also demonstrated that validating full SHACL features is an NP-hard problem. Building on these insights, they proposed specific fragments of SHACL that are computationally tractable, along with a fundamental algorithm for validating shape schemas using SPARQL [39]. In a related vein, Andresel et al. [8] propose a stricter semantics for recursive SHACL, drawing inspiration from stable models employed in Answer Set Programming (ASP). This innovative approach enables the representation of SHACL constraints as logic programs and leverages existing ASP solvers for shape schema validation. Importantly, this approach allows for the inclusion of negations in recursive validations. Further, Figuera et al. [56] present Trav-SHACL, an approach that focuses on query optimization techniques aimed at enhancing the incremental behavior and scalability of shape schema validation.
@@
While SHACL has been adopted in a broad range of use cases, given a large graph it remains a challenge how to define shapes efficiently [130]. In many industrial settings with billions of entities and facts [119] creating shapes manually simply is not an option. The current state of the art can automatically extract shapes on WikiData (ca. 2 Billion facts) in less than 1.5 hours while filtering shapes based on the well-established notions of support and confidence to avoid reporting thousands of shapes that are so rare or apply to such a small subset of the data that they become meaningless [131]. Still, more work is needed to increase scalability further and also to help users make good use of the mined shapes [132] and, e. g., interactively use them to correct and improve the quality of their graphs.

### v2

@@ 4.2 Ontologies
An ontology is commonly defined as a formal, machine-readable representation of key concepts and relationships within a specific domain [GuarinoOberleStaabWhatIsAnOntology2009, OberleMiddleware2006]. In essence, ontologies capture a shared perspective [GuarinoOberleStaabWhatIsAnOntology2009] that is, the formal conceptualization of ontologies expresses a consensus view among different stakeholders. Visualizing ontologies is akin to viewing a spectrum, with a specificity of concepts, their relationships, and the granularity of meaning representation varying along this continuum [DBLP:conf/dagstuhl/McGuinness03, DBLP:journals/sigmod/UscholdG04, DBLP:journals/ao/Uschold15]. A controlled vocabulary corresponds to the less expressive form of ontology, comprising a restrictive list of words or terms used for labeling, indexing, or categorization. The Clinical Data Interchange Standards Consortium (CDISC) Terminology is an exemplary vocabulary that harmonizes definitions in clinical research. [^42] https://datascience.cancer.gov/resources/cancer-vocabulary/cdisc-terminology A thesaurus is located next in the spectrum; they enhance controlled vocabularies with information about terms and their synonyms and broader/narrower relationships. The Unified Medical Language System (UMLS) integrates medical terms and their synonyms. [^43] https://www.nlm.nih.gov/research/umls/index.html Next, taxonomies are built over controlled vocabularies to provide a hierarchical structure, e. g., parent/child relationship. SNOMED-CT [^44] https://www.snomed.org/value-of-snomedct (Systematized Nomenclature of Medicine Clinical Terms) provides a terminology and coding system used in healthcare and medical fields; medical concepts organized in a hierarchical structure enabling a granular representation of clinical information. The Simple Knowledge Organization System (SKOS) [^45] https://www.w3.org/TR/skos-reference/ is a W3C standard to describe knowledge about organizational systems. Lastly, ontologies are at the highest extreme of the spectrum, integrating sets of concepts with attributes and relationships to define a domain of knowledge.
Note, SKOS is a popular standard for modeling domain-specific taxonomies in the different scientific communities such as economics, social sciences, etc. to represent concepts and their relationships, most importantly narrower, broader, and related. However, it does not have the expressiveness of OWL with its complex expressions on classes and relations. For a detailed discussion, we refer to the literature such as [DBLP:conf/owled/JuppBS08] and the W3C on using OWL and SKOS [^46] https://www.w3.org/2006/07/SWD/SKOS/skos-and-owl/master.html.
@@ 4.3 Types and Examples of Ontologies
A network of ontologies, such as the example shown in Figure 1, may consist of a variety of ontologies created by different actors and communities. Ontologies may be the result of a transformation or reengineering activity of a legacy system, such as a relational database or existing taxonomy such as the Dewey Decimal Classification [^47] https://www.oclc.org/en/dewey.html or Dublin Core. Other ontologies are created from scratch. This involves applying existing methods and tools for ontology engineering and choosing an appropriate representation language for the ontology (see Section 6).
@@
Ontology engineering deals with the methods for creating ontologies [GomezPerezFernandezLopezOntologicalEngineering] and has its origins in software engineering in the creation of domain models and in database design in the creation of conceptual models. A good overview of ontology engineering can be found in several reference books [GomezPerezFernandezLopezOntologicalEngineering]. Ontologies vary greatly in their structure, size, development methods applied, and application domains considered. Complex ontologies are also distinguished in terms of their purpose and granularity.
@@
Domain Ontologies represent knowledge specific to a particular domain [Euzenat2007a, OberleMiddleware2006]. Domain ontologies are used as external sources of background knowledge [Euzenat2007a]. They can be built on foundational ontologies [Oberle:2007:DES:1290204.1290385] or core ontologies [ScherpEtAlDesigningCoreOntologiesIOS2011], which provide precise structuring to the domain ontology and thus improve interoperability between different domain ontologies. Domain ontologies can be simple such as the FOAF ontology or the event ontology mentioned above, or very complex and extensive, having been developed by domain experts, such as the SNOMED medical ontology.
@@
Core Ontologies represent a precise definition of structured knowledge in a particular domain spanning multiple application domains [ScherpEtAlDesigningCoreOntologiesIOS2011, OberleMiddleware2006]. Examples of core ontologies include core ontologies for software components and web services [OberleMiddleware2006], for events and event relationships [emf-mtap], or for multimedia metadata [www-m3o]. Core ontologies should thereby build on foundational ontologies to benefit from their formalization and strong axiomatization [ScherpEtAlDesigningCoreOntologiesIOS2011]. For this purpose, new concepts and relations are added to core ontologies for the application domain under consideration and are specialized by foundational ontologies.
@@
Foundational Ontologies have a very wide scope and can be reused in a wide variety of modeling scenarios [BorgoMasoloDOLCE2009]. They are therefore used for reference purposes [OberleMiddleware2006] and aim to model the most general and generic concepts and relations that can be used to describe almost any aspect of our world [BorgoMasoloDOLCE2009, OberleMiddleware2006], such as objects and events. An example is the Descriptive Ontology for Linguistic and Cognitive Engineering (DOLCE) [BorgoMasoloDOLCE2009]. Such basic ontologies have a rich axiomatization that is important at the developmental stage of ontologies. They help ontology engineers to have a formal and internally consistent conceptualization of the world, which can be modeled and checked for consistency. For the use of foundational ontologies in a concrete application, i. e., during the runtime of an application, the rich axiomatization can often be removed and replaced by a more lightweight version of the foundational ontology.
@@ 4.4 Distributed Network of Ontologies and Ontology Patterns
A network of ontologies must be flexible with respect to the functional requirements imposed on it. This is because systems are modified, extended, combined, or integrated over time. In addition, the networked ontologies must lead to a common understanding of the modeled domain. This common understanding can be achieved through a sufficient level of formalization and axiomatization, and through the use of ontology patterns. An ontology pattern, similar to a design pattern in software engineering, represents a generic solution to a recurring modeling problem [ScherpEtAlDesigningCoreOntologiesIOS2011]. Ontology patterns allow to select parts from the original ontology. Either all or only certain patterns of an ontology can be reused in the network. Thus, to create a network of ontologies, e. g., existing ontologies and ontology patterns can be merged on the Web. The ontology engineer can drive or explicitly provide for the modularization of ontologies using ontology patterns. Core ontologies represent one approach to designing a network of ontologies (see in detail [ScherpEtAlDesigningCoreOntologiesIOS2011]). They allow to capture and exchange structured knowledge in complex domains. Well-defined core ontologies fulfill the properties mentioned in the previous section and allow easy integration and smooth interaction of ontologies (see also [ScherpEtAlDesigningCoreOntologiesIOS2011]). The networked ontologies approach leads to a flat structure, as shown in Figure 1, where all ontologies are used on the same level. Such structures can be managed up to a certain level of complexity.
@@
The approach of networked core ontologies is illustrated by the example of ontology layers starting from foundational to core to domain ontologies. As shown in Figure 4, DOLCE is the foundational ontology at the bottom layer, the Multimedia Metadata Ontology (M3O) [www-m3o] as the core ontology for multimedia metadata, and an extension of M3O for the music domain. Core ontologies are typically defined in description logic and cover a field larger than the specific application domain requires [GangemiPresuttiODP2009]. Concrete information systems will typically use only a subset of core ontologies. To achieve modularization of core ontologies, they should be designed using ontology patterns. By precisely matching the concepts in the core ontology with the concepts provided in the foundational ontology, they provide a solid foundation for future extensions. New patterns can be added and existing patterns can be extended by specializing the concepts and roles. Figure 4 shows different patterns of the M3O and DOLCE ontologies.
Ideally, the ontology patterns of the core ontologies are reused in the domain ontologies [GangemiPresuttiODP2009], as shown in Figure 4. However, since it cannot be assumed that all domain ontologies are aligned with a foundational or core ontology, the option that domain ontologies are developed and maintained independently must also be considered. In this case, domain knowledge can be reused in core ontologies by applying the Descriptions and Situations (DnS) ontology pattern of the foundational ontology DOLCE. The DnS ontology pattern is an ontological formalization of context [OberleMiddleware2006] by defining different views using roles. These roles can refer to domain ontologies and allow a clear separation of the structured knowledge of the core ontology and domain-specific knowledge. To model a network of ontologies, such as the example described above, the Web Ontology Language (OWL) and its ability to axiomatize using description logic [DBLP:conf/dlog/2003handbook] is used. In addition to being used to model a distributed knowledge representation and integration, OWL, is also used in particular to derive inferences from this knowledge, which is described in Section 6.
@@ 5.1 Graph Data Creation
Graph data can be created by transforming legacy data via a data integration system [lenzerini2002data], which consists of a unified schema, data sources, and mapping rules. These mapping rules define the concepts within the schema and establish links to the data sources. By employing declarative definitions, knowledge graph creation promotes modularity and reusability. This approach allows users to trace the entire graph creation process, leading to improved transparency and ease of maintenance.
@@
The mapping language R2RML [W3CR2RML] defines mapping rules from relational databases (relational data models) to RDF graphs. These mappings themselves are also RDF triples [W3TURTLE]. Because of its compact representation, Turtle is considered a user-friendly notation of RDF graphs. The structure of R2RML is illustrated in Figure 5; essentially, table contents are mapped to triples by the classes SubjectMap, PredicateMap, and ObjectMap. If the object is a reference to another table, this reference is called RefObjectMap. Here, SubjectMap contains primary key attributes of the corresponding table. Thus, there exists a mapping rule representable in RDF graphs by means of which tables of relational databases can be represented as RDF graphs.
Figure 5: Structure of a relational data mapping (source: [W3CR2RML]).
@@
The RDF Mapping Language (RML)[DimouSCVMW14] extends R2RML to encompass the definition of logical sources in various formats, including CSV, JSON, XML, and HTML. This enhancement enables RML to introduce new operators that facilitate the integration of data from diverse sources into the Semantic Web. Thus, instead of LogicalTable, RML includes the tag LogicalSource, to allow for the retrieval of data in several formats. Additionally, RML resorts to W3C-standardized vocabularies and enables the definition of retrieval procedures to collect data from Web APIs or databases. R2RML and RDF mapping rules are expressed in RDF, and their graphs document how classes and properties in one or various ontologies that are part of an RDF graph are populated from data collected from potentially heterogeneous data sources.
@@
Over time, the Semantic Web community has actively contributed to addressing the challenge of integrating heterogeneous datasets, resulting in the development of several frameworks for executing declarative mapping rules [chebotko2009semantics, calvanese2017ontop, DBLP:conf/www/PriyatnaCS14]. A rich spectrum of tools (e. g., RMLMapper [DimouSCVMW14], RocketRML [csimcsek2019rocketrml], CARML [^48] https://github.com/carml/carml, SDM-RDFizer [iglesias2020sdm], Morph-KGC [arenas2022morph], and RMLStreamer [oo2022rmlstreamer]) offers the possibility of executing R2RML and RML rules and efficiently materializing the transformed data into RDF graphs. Van Assche et al. [DBLP:journals/ws/AsscheDHHMD23] provided an extensive survey detailing the main characteristics of these engines. Despite significant efforts in developing these solutions, certain parameters can impact the performance of the graph creation process [Chaves-FragaEIC19]. Existing engines may face challenges when handling complex mapping rules or large data sources. Nonetheless, the community continues to collaborate and address these issues. An example of such collaboration is the Knowledge Graph Construction Workshop 2023 Challenge [^49] https://zenodo.org/record/7689310 that took place at ESWC 2023. This community event aims to understand the strengths and weaknesses of existing approaches and devise effective methods to overcome existing limitations.
@@
RDF graphs can also be dynamically created through the execution of queries over data sources. These queries involve the rewriting of queries expressed in terms of an ontology, based on mapping rules that establish correspondences between data sources and the ontology. Tools such as Ontop [calvanese2017ontop], Ultrawrap [DBLP:journals/ws/SequedaM13], Morph [DBLP:conf/www/PriyatnaCS14], Squerall [mami2019squerall], and Morph-CSV [chaves2020enhancing] exemplify systems that facilitate the virtual creation of RDF graphs.
@@ 5.2 Quality and Validation of Graph Data
Quality and validation of the graph data are crucial to maintaining the integrity of the Semantic Web [DBLP:journals/corr/abs-2308-14217, DBLP:journals/semweb/Debattista0AC18, Zaveri2015surveyQuality]. The evaluation of integrity constraints allows for the identification of inconsistencies, inaccuracies, or contradictions within the data. They also help maintain consistency by ensuring related data elements remain coherent. Constraints are logical statements – expressed in a particular language – that impose restrictions on the values taken for target nodes in a given property.
Constraints can be expressed using OWL [DBLP:conf/aaai/TaoSBM10], SPARQL queries [DBLP:conf/edbt/LausenMS08], or using shapes. However, the interpretation of the results depends on the semantics followed to interpret the failure of an integrity constraint. For example, constraints expressed in OWL are validated using an Open-World Assumption (OWA) (i. e., a statement cannot be inferred to be false based on failures to prove it) and under the absence of the Unique Name Assumption (UNA) (i. e., two different names may refer to the same object). These two features make it difficult to validate data in applications where data is supposed to be complete. Definitions of integrity constraint semantics in OWL using the Closed-World Assumption [Motik2007, Motik2009, DBLP:conf/aaai/TaoSBM10] overcome these issues.
@@
Data quality conditions and integrity constraints can also be expressed as graphs of shapes or the so-called shapes schema. A shape corresponds to a conjunction of constraints that a set of nodes in an RDF graph must satisfy [DBLP:series/synthesis/2021Hogan]. These constraints can restrict the types of nodes, the cardinality of certain properties, and the expected data types or values for specific properties. A shape can target the instances of a class, the domain or range of a property, or a specific node in the RDF graph. A shape or node in a shapes graph is validated in an RDF graph, if and only if, all the target nodes in the RDF graph satisfy all the constraints in the shape. Figure 6 presents a shapes graph of three shapes targeting the classes Brand, Playcount, and MusicArtist. Each of the shapes comprises one constraint. In the shapes Brand and MusicArtist the properties title and name can take more than one value, while the shape Playcount states that each instance of the class Playcount must have exactly one value of the property count. Additionally, the instances of the class Brand must be related to valid instances of the class Playcount which should also be related to valid instances of the class MusicArtist.
@@
There are two standards for defining shapes, ShEx (Shape Expressions) [SHEX]) and SHACL (Shapes Constraint Language) [SHACL]). Both define shapes over the attributes (i. e., owl:DatatypeProperties), and constraints on incoming/outgoing arcs, cardinalities, RDF syntax, and extension mechanism. These inter-class constraints induce a shape network used to validate the integrity and data quality properties of an RDF graph.
@@
SHACL and ShEx, although sharing a common goal, adopt distinct approaches. ShEx seeks to offer a language serving as a grammar or schema for RDF graphs, delineating RDF graph structures for validation. On the other hand, SHACL is positioned as the W3C recommendation for validating RDF graphs against a conjunction of constraints, emphasizing a constraint language for RDF. Despite their analogous roles in specifying shapes and constraints for RDF data, ShEx and SHACL differ in syntax, expressiveness, and community adoption [DBLP:series/synthesis/2017Gayo].
@@
The evaluation results of a SHACL shape network over an RDF graph are presented in validation reports using a controlled vocabulary. A validation report includes explanations about the violations, the severity of the violation, and a message describing the violation. SHACL is the language selected by the International Data Space (IDS) to express the restrictions that state the integrity over RDF graphs [DBLP:books/sp/22/0002LB22]. Besides the integrity validation of an RDF graph, SHACL can be utilized to describe data sources and the certification of a query answer [DBLP:conf/vldb/Rohde21], as metadata to enhance the performance of a SPARQL query engine [DBLP:conf/edbt/RabbaniLH21], to certify access policies [DBLP:conf/esws/RohdeV23], and to provide provenance as a result of the validation of integrity constraints [DBLP:conf/edbt/DelvaDJB23].
@@
In the context of a quality assessment pipeline, one crucial step involves validating the shape schema against a graph. It is important to mention that the validation of recursive shape schemas is not explicitly addressed in the SHACL specification [SHACL]. To address this gap, Corman et al. [Corman2018] introduce a semantic framework for validating recursive SHACL. They also demonstrated that validating full SHACL features is an NP-hard problem. Building on these insights, they proposed specific fragments of SHACL that are computationally tractable, along with a fundamental algorithm for validating shape schemas using SPARQL [Corman2019]. In a related vein, Andresel et al. [Andresel2020] propose a stricter semantics for recursive SHACL, drawing inspiration from stable models employed in Answer Set Programming (ASP). This innovative approach enables the representation of SHACL constraints as logic programs and leverages existing ASP solvers for shape schema validation. Importantly, this approach allows for the inclusion of negations in recursive validations. Further, Figuera et al. [DBLP:conf/www/FigueraRV21] present Trav-SHACL, an approach that focuses on query optimization techniques aimed at enhancing the incremental behavior and scalability of shape schema validation.
@@
While SHACL has been adopted in a broad range of use cases, given a large graph it remains a challenge how to define shapes efficiently [DBLP:conf/www/RabbaniLH22]. In many industrial settings with billions of entities and facts [Noy19] creating shapes manually simply is not an option. The current state of the art can automatically extract shapes on WikiData (ca. 2 Billion facts) in less than 1.5 hours while filtering shapes based on the well-established notions of support and confidence to avoid reporting thousands of shapes that are so rare or apply to such a small subset of the data that they become meaningless [DBLP:journals/pvldb/RabbaniLH23]. Still, more work is needed to increase scalability further and also to help users make good use of the mined shapes [DBLP:conf/sigmod/RabbaniLH23] and, e. g., interactively use them to correct and improve the quality of their graphs.